7. The body, weighing 1kg, slides across the sloping plane with length l = 2m and height h = 0.4m Find the work done by each force acting on the body as well as the general work, when the body traverses the sloping plane: a) without friction (u = 0); b by friction (u = 0.04).


1
Expert's answer
2021-04-01T18:41:51-0400

a)


"W_G=mgh=(1)(9.8)(0.4)=3.9\ J\\W=3.9\ J"

b)


"W_G=mgh=(1)(9.8)(0.4)=3.9\ J\\W_F=\mu mgl\cos{\alpha}\\W_F=0.04(1)(9.8)(2)\frac{\sqrt{2^2-0.4^2}}{2}=0.8\ J\\W=3.9-0.8=3.1\ J"